Please help me solve this problem Solve by Addition method x-y+z=4
2x=y-3z=-5
x+3y+2z=2

Answer by Edwin McCravy(20055)


 x -  y +  z =  4
2x +  y - 3z = -5
 x + 3y + 2z =  2

Choose a letter to eliminate, and
choose two of the equations that contain
that letter to eliminate it from.

I choose to eliminate y from the first two
equations since all I have to do is add
the equations as they are and the y's
will cancel out

 x -  y +  z =  4
2x +  y - 3z = -5
-----------------
3x      - 2z = -1

Choose the equation you haven't chosen yet,
and another one that you have already used,
to eliminate the SAME letter you eliminated
before.

I haven't used the 3rd one, so I will choose
it and the first one to eliminate the SAME
letter I eliminated before, which is y:

 x -  y +  z =  4
 x + 3y + 2z =  2

To eliminate y we must first multiply the
first equation by 3:

3x - 3y + 3z = 12
 x + 3y + 2z =  2
-----------------
4x      + 5z = 14

So we take the two equations with y eliminated
and solve them:

3x  - 2z = -1
4x  + 5z = 14

Now we eliminate z by multiplying the first by 5
and the second by 2

15x  - 10z = -5
 8x  + 10z = 28
---------------
23x        = 23
   
       23x = 23
         x = 1

Substitute 1 for x in

 3x  - 2z = -1
3(1) - 2z = -1
   3 - 2z = -1
      -2z = -4
        z = 2

Substitute x = 1 and z = 2 into
one of the original equations,
say the first:

 x -  y + z =  4
 1 -  y + 2 =  4
     -y + 3 =  4 
         -y =  1
          y = -1

Solution: (x, y, z) = (1, -1, 2)

Edwin